When saving = Investment, show that the trade balance equals the government spendig balance



1
Expert's answer
2022-02-07T08:42:06-0500

Macroeconomics

When saving is equal to investment, we apply the fundamental macroeconomic definition, savings are the total income less the expenditures, while investments are physical investments.

People on the other hand, only from consumption and purchases made by government. the remainder of the 3 sector mode is investment, which is categorized as savings.

"Y=C+I+G+(X-M)"

"Y = Total\ real\ income"

"C+I+G+(X-M)=Total\ goods\ produced"

"Total\ real\ incom=Total\ goods\ produced"

"C+G= Consumption\ by\ people"

"I= Savings"

Trade balance is the balance in trade, resulting from the difference between exports and imports.

Government spending balance, is balance is arrived at from subtracting expenditures of the government, from its total revenue.

"Government\ Revenue=(X-M)" "= \bold{Trade\ balance}"

"Government\ Expenditure =M"

"Government\ Revenue=X"

"Government\ spending\ balance= (X-M




)"

"\therefore \bold{Trade\ balance=Government\ spending\ balance.}"

Reference.

Lepenies, P. (2019). Products before People: How Inequality Was Sidelined by Gross National Product. InΒ Histories of Global InequalityΒ (pp. 83-105). Palgrave Macmillan, Cham. https://doi.org/10.1007/978-3-030-19163-4_4